# Judging LLMs on a Simplex

### v1

@@ 3.1 Some basic geometry
Assumption 2.
Judge $\hat{s}_{j}$ satisfies “moderate constancy” if its confusion matrix is the same for all non-self candidates: For each $m$, there is some $\theta^{(j)}_{m}$ such that $\theta^{(j)}_{m,k}=\theta^{(j)}_{m}$ for all $k\neq j$ [^1] For simplicity, the theoretical results study the case where self-judging is defined as the same exact LLM is used to judge itself. In practice, we adjust for self-judging by disallowing each LLM judge from judging its family (e.g., Claude 3.5 Haiku is not allowed to judge Claude 3.5 Sonnet and Claude 3.7.)
Under such constancy assumptions, we can jointly visualize all candidates that are consistently scored by judge $j$ as points within the convex hull of the judge’s vertices. (For notational ease, we omit the judge superscript when discussing individual judges.) The candidates’ barycentric coordinates become sufficient for determining their relative performance. For instance, under strong constancy in a 2-level scoring system, every candidate $k$ falls between $\theta_{1}$ and $\theta_{2}$, where its barycentric coordinates (as well as its true mean score) are determined by the relative distance of $\gamma_{k}$ to vertices $\theta_{1}$ and $\theta_{2}$ (Fig 3.2 left). Similarly, under strong constancy in a 3-level scoring system, every candidate $k$ falls in the triangle $\triangle_{\theta_{1},\theta_{2},\theta_{3}}$ (Fig 3.2 middle) [^2] Note that the probability of score 1 is implicit since probabilities sum to 1, allowing the 2-D simplex to be projected onto a triangle. and its barycentric coordinates are determined by relative area of its subtriangles (Fig 3.3 c), i.e.,
$\pi_{k,1}=\frac{\text{Area of }\triangle_{(\gamma_{k},\theta_{2},\theta_{3})}}{\text{Area of }\triangle_{(\theta_{1},\theta_{2},\theta_{3})}},\quad\pi_{k,2}=\frac{\text{Area of }\triangle_{(\gamma_{k},\theta_{1},\theta_{3})}}{\text{Area of }\triangle_{(\theta_{1},\theta_{2},\theta_{3})}},\quad\pi_{k,3}=\frac{\text{Area of }\triangle_{(\gamma_{k},\theta_{1},\theta_{2})}}{\text{Area of }\triangle_{(\theta_{1},\theta_{2},\theta_{3})}}.$
So comparing true score prevalences reduces to comparing triangle areas.
Furthermore, because $\left(\begin{matrix}\gamma_{k}\\ E[S^{*}_{k}]\end{matrix}\right)=\sum_{m=1}^{M}\pi_{k,m}\left(\begin{matrix}\theta_{m,k}\\ m\end{matrix}\right),$ we can visualize each candidate’s true score by augmenting each judge vertex with the true score, drawing the triangle between these augmented vertices, and finding the intersection between the triangle and a vertical line at each candidate (Fig 3.2 right) Then the true score of each candidate corresponds to the height of this intersection. That is, comparing performance of candidates reduces to comparing heights.
We note that both constancy assumptions are weaker than much of the existing literature. For instance, if one were to average judge-assigned scores directly, the implicit assumption is that LLM scores are unbiased, i.e., $\mathbb{E}[\hat{S}_{k}]=\mathbb{E}[\hat{S}_{k}^{*}]$. This assumes the judge vertices fall in very specific locations.
Using these geometric properties, we can build intuition for how candidate rankings can be determined. For instance, if the judge vertices are known, we can determine the barycentric coordinates for every candidate as long as every candidate is consistently scored by at least one judge, which then implies that we can recover the true scores for all candidates. To understand the theoretical limits of identifiability when gold-standard labels are unavailable, we examine identifiability when the judge vertices are unknown. For the following results, we present only the geometric intuition; detailed proofs are in the Appendix. We begin with 2-level scoring systems and then build up to 3+ levels.
3.2 2-level scoring systems
When access to the gold-standard scoring function $s^{*}$ is unavailable, the location of judge vertices generally cannot be recovered even with access to infinitely large benchmark datasets, as $\theta_{1}$ and $\theta_{2}$ can be any two points along the segment $(0,1)$ to $(1,0)$ that envelopes the candidates (Figure 3.3 left). So in general, neither constancy assumption is sufficient for recovering the absolute values of the candidate scores $\mathbb{E}[S^{*}_{k}]$, even if one increased the number of judges. [^3] There is an edge case where we can recover rankings under the constancy assumption. If we have at least one candidate with only judge-assigned scores of 1s and another candidate with only 0s, the constancy assumption implies that the judge is perfect and the true score coincides with the judge-assigned score.
Figure 3.3: When gold-standard labels are unavailable, the judge performance characteristics, and thus judge vertices, are unknown. Multiple configurations are possible given candidates $k$=1 and $k$=2. Two such configurations ($𝜃\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}{\theta}$ vs $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}{\theta^{\prime}}$) are shown for 2-level scoring systems (a,b) and 3-level scoring systems (c,d); a 3-D version is shown in (e). While prevalence and score rankings are the same under different configurations for the 2-level scoring system, these rankings are flipped between the two configurations for the 3-level scoring system.
Nevertheless, under strong constancy, we can see that the position of each candidate along the line is sufficiently informative for the purposes of ranking. Assuming the judge doesn’t prefer to assign score 2 when the true score is 1 (known as the “label-flipping problem” [38]), a candidate with higher prevalence of judge-assigned scores of 2 is indeed better. Thus we can rank candidates by their relative positions along the line (or equivalently, their relative barycentric coordinates) as long as the judges satisfy a weak monotonicity condition:
Assumption 3.
@@
The $j$-th judge’s probability of assigning the lowest score when the true score is equal to $m$ is decreases with respect to $m$.
Under moderate constancy, we use a similar idea. For candidates that can only be evaluated by a subset of judges, we can determine their position relative to candidates that can be evaluated by all judges, which can then be used to rank all candidates. This leads to the following positive result:
@@
For 2-level scoring systems, the rank of candidates is identifiable from the distribution of assigned scores if: (i) there is at least one judge that satisfies Assumptions 1 and 3 or (ii) there are at least two judges that satisfy Assumptions 1 and 3 and at least four candidates.
@@
The story is more complex for scoring systems with 3 or more levels. To build intuition, we begin with asking whether the score prevalences can be ranked, and then whether the true candidate scores can be ranked. We discuss the geometric intuition for 3-level scoring systems below, but the same arguments can be readily extended to more levels.
(i) Ranking prevalences. WLOG, suppose we wanted to rank the prevalence of the true score $m=3$. Under strong constancy, the denominators for determining the barycentric coordinates are the same, so we simply need to compare the areas of the subtriangles $\triangle_{(\gamma_{k},\theta_{1},\theta_{2})}$ (Fig 3.3 c). This boils down to comparing the distances from $\gamma_{k}$’s to the line $\overrightarrow{\theta_{1}\theta_{2}}$, which can be determined solely from the slope of this line (not the exact locations of $\theta_{1}$ and $\theta_{2}$). However, in the absence of gold-standard labels, the slope of $\overrightarrow{\theta_{1}\theta_{2}}$ is generally unknown, as there are an infinite number of triangles that envelope the candidates (Fig 3.3 c vs d). So neither absolute nor relative prevalences can be determined, in general.
While this negative result may be disappointing, this analysis also shows that there is hope: if one had strong prior beliefs about the slope, one can determine which prevalence rankings are more likely. In addition, if candidates have extremely different distributions of judge-assigned score, the differences may be so large that the ranking is undebatable, regardless of one’s prior beliefs. Similarly, if an LLM judge satisfies the constancy assumption for a wide variety of candidates, the number of possible triangles enveloping all candidates, and thus the possible prevalence rankings, becomes very limited (see Appendix for visualizations).
(ii) Ranking mean scores. Given the nonidentifiability of prevalence rankings, it is unsurprising that the ranking of true candidate scores is also nonidentifiable when gold standard labels are unavailable. To see this directly, note that the relative scores, which corresponds to the relative heights of each candidate’s vertical projection onto the augmented triangle (Fig 3.2 right), varies depending on where the judge’s vertices are located. If a judge’s vertices are shifted enough, we may flip the relative heights between candidates; this is certainly possible if candidates are closer together (Fig 3.3 e). This argument applies to both constancy assumptions and any number of judges. While this negative result again may be disappointing, there are various ways to filter down possible rankings as outlined above.
These results highlight that ranking uncertainty comes not just from sampling variation (aleatoric uncertainty) but also uncertainty about which assumptions hold (epistemic uncertainty). Putting these results together, we have the following.
@@
Consider grading questions across 3+ levels and suppose one has $J$ judges that satisfy Assumptions 1 and 3 or Assumptions 1 and 3. Given only the distribution of judge-assigned scores, there exist candidates whose prevalence and score rankings cannot be identified.
4 Bayesian Ranking on a Simplex
The identifiability analyses in Section 3 reveal the significant role of epistemic uncertainty. For binary scoring systems, ranks can be identified under relatively mild constancy and monotonicity assumptions; for 3+-level scoring systems, ranks are generally non-identifiable unless much stronger prior information is provided. As such, we use a Bayesian framework to explicitly control the assumptions and prior information used for ranking estimation and inference. In particular, we use the Bayesian prior to both take away information by exploring relaxations of the constancy assumption and injecting information by specifying likely slopes between judge vertices. We present these key pieces of the Bayesian model below. Additional implementation details and recommendations for priors are provided in the Appendix.
@@ 4.1 Probability model
We model the assigned score $\hat{S}_{ik}^{(j)}$ by judge $j$ to candidate $k$’s answer to the $i$-th question, given its true score $S_{ik}^{*}$, as independent draws from a multinomial distribution with parameter $\theta_{S_{ik}^{*},k}^{(j)}$. Marginalizing out the true latent scores, the likelihood of the observed data is
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big{[}\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big{]},$ (4.2)

### v2

@@ 4 Theoretical limits of ranking identifiability
Judge $j$’s probability of assigning the lowest score when the true score is equal to $m$ decreases with respect to $m$.
The story under Moderate Constancy is similar. Using similar geometric reasoning (see Appendix), we find that the only difference is that at least two judges and four candidates are needed to obtain enough information to recover the true ranking.
@@
For 2-level scoring, any of the following conditions are sufficient for candidate rankings to be identifiable from the distribution of judge-assigned scores: (i) There is $J=1$ judge and it satisfies Assumptions 4 and 3. (ii) There are $K\geq 4$ candidates, $J\geq 2$ judges, and the judges satisfy Assumptions 5 and 3.
Remark. For completeness, we present in the Appendix necessary and sufficient conditions for recovering rankings in two-level systems. The geometry is complex, so we defer the details to the appendix.
@@
The story is more nuanced for scoring systems with 3 or more levels. Below, we present geometric arguments for 3-level scoring systems, but the same arguments extend to more levels. To build intuition, we first discuss the tasks of (i) ranking the prevalence of each score and then (ii) ranking the expected scores.
(i) Ranking prevalences. Without loss of generality, suppose we wanted to rank the prevalence of true score $m=1$. As mentioned in Section 3, this is equivalent to ranking the barycentric coordinates $\pi_{k,1}$ across candidates $k=1,\cdots,K$, which are defined as the area ratios (1). Under strong constancy, the denominators of these ratios are the same for all $k$, so ranking $\pi_{k,1}$ reduces to ranking the areas of the subtriangles $\triangle_{(\gamma_{k},\theta_{2},\theta_{3})}$ Nevertheless, we find that the rankings are non-identifiable in general, even under strong constancy. Figure 4 provides a visual proof: panels (a) and (b) show two equally valid judge configurations ($\{\theta_{1},\theta_{2},\theta_{3}\}$ versus $\{\theta^{\prime}_{1},\theta^{\prime}_{2},\theta^{\prime}_{3}\}$) that can explain candidate score distributions $\gamma_{k=1}$ and $\gamma_{k=2}$, as both judge configurations envelop the candidate points. However, the areas flip between configurations. In panel (a), subtriangle $\triangle_{(\gamma_{1},\theta_{2},\theta_{3})}$ has greater area than $\triangle_{(\gamma_{2},\theta_{2},\theta_{3})}$, while in panel (b) this relationship reverses. Thus the two panels show that even under strong constancy, the rankings of the score prevalences cannot be recovered if there is no prior knowledge about judge quality.
Figure 4: Non-identifiability in 3-level scoring. Same candidate positions (green/red) explained by different judge configurations yield opposite prevalence rankings: $\pi_{1,1}>\pi_{2,1}$ in (a) vs. $\pi_{1,1}<\pi_{2,1}$ in (b).
(ii) Ranking expected scores. Given the nonidentifiability of prevalence rankings, it is unsurprising that the ranking of true candidate scores is also nonidentifiable when gold standard labels are unavailable. To see this more directly, notice that the expected scores of the candidates, which correspond to the relative heights of each candidate’s vertical projection onto the augmented simplex (Figure 2c), varies based on the location of the judge vertices. If a judge’s vertices are shifted enough, the height rankings between candidates can shift. This argument applies, under either constancy assumption and for any number of judges.
Together, these results illustrate that epistemic uncertainty can significantly affect recoverability of the true rankings, even when aleatoric uncertainty is removed.
@@
Consider the 3+-level scoring setting with all judges either satisfying Assumptions 4 and 3 or Assumptions 5 and 3. Given only the distribution of judge-assigned scores, there exist candidates whose prevalence and/or expected-score rankings cannot be identified.
4.3 Recoverability is dataset-specific
The prior theoretical results show that recoverability depends not only on judge quality assumptions but also the problem setting. In particular, Theorems 1 and 2 uncover a dependence on the number of scoring levels, supporting the “folk wisdom” that LLM judges tend to be more effective for 2-level scoring systems and less effective for 3+levels (Shankar2024-an; Husain2025-js). Leveraging the geometric perspective further, we find that recoverability is, in fact, dataset-specific, as illustrated in the following two examples. This means we need a way to rigorously quantify how sensitive the estimated rankings for a given dataset are to varying levels of epistemic uncertainty, which we address using Bayesian inference in the next section.
Example 1: Dataset requiring stronger priors. Consider again the Figure 4 example and the task of ranking prevalence of score $m=1$. Intuitively, the judge configuration in panel (a) may seem more probable than panel (b) if we believe that LLM judges can distinguish scores 2 and 3 moderately well. To formalize this prior belief, we distill it using geometric arguments. In particular, ranking $\pi_{k,1}$ is equivalent to ranking areas of subtriangles $\triangle{\gamma_{k},\theta_{2},\theta_{3}}$. Because all the subtriangles have the same line segment $\overline{\theta_{2}\theta_{3}}$ as the base, ranking areas is equivalent to ranking distances from each candidate $\gamma_{k}$ to the line $\overrightarrow{\theta_{2}\theta_{3}}$. Because ranking distances only require knowing the slope of this line, it is sufficient to state our prior belief about judge quality in terms of slopes. This is something we generally have a good prior about. In contrast, it is generally more difficult to specify one’s prior about the exact location of the judge vertices, since LLMs are often miscalibrated (Wang2024-gp).
Example 2: Dataset allowing weaker priors. Consider a candidate receiving mostly 1’s versus another receiving mostly 3’s from an LLM judge (see visualization in Appendix). When score distributions are so different, the ranking between two candidates may be quite evident, even if we have limited prior knowledge on the performance of the LLM judges and even if the constancy assumptions do not hold. However, we need a method to quantify when constancy violations flip rankings, as discussed in the next section.
5 Bayesian framework
Our theoretical findings and geometric framework have shown us that how much epistemic uncertainty impacts our ability to rank LLM candidates is dataset-specific. To quantify this more formally, we have designed a Bayesian framework that encodes epistemic uncertainty in terms of two priors: (i) how strongly the constancy assumption holds, as represented through random effects (Section 5.2), and (ii) prior beliefs about judge discriminative ability, as represented by slopes between judge vertices (Section 5.3). We can then conduct sensitivity analyses that assess how robust the estimated rankings are to varying levels of epistemic uncertainty by systematically varying the strength of the Bayesian priors. Moreover, by defining appropriate hyperpriors, Bayesian posterior inference can fully incorporate our epistemic uncertainty into the posterior distribution. Due to limited space, we highlight only the key pieces of the Bayesian framework here and leave the full specification to the Appendix.
@@ 5.1 Probability model
We model the assigned score $\hat{S}_{ik}^{(j)}$ by judge $j$ to candidate $k$’s answer to the $i$-th question, given its true score $S_{ik}^{*}$, as independent draws from a multinomial distribution with parameter $\theta_{S_{ik}^{*},k}^{(j)}$. [^2] Note that the model deliberately ignores the correlation between judge scores to avoid potential misspecification of complex correlation structures; we prove in the Appendix that the posterior inference procedure is still consistent. Marginalizing out the true latent scores, the likelihood of the observed data is
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big],$
where $n$ is the number of questions. To remove the influence of self-preference, note that we filter for $k\neq j$.
5.2 Prior for relaxing constancy
To relax the constancy assumptions (Assumptions 4 and 5), we model the prevalence of each judge-assigned score for a given candidate using random effects relative to a “base” score prevalence $\vec{\pi}_{k}$:
$\vec{\pi}_{k}^{(j)}=(1-W_{k}R_{j})\vec{\pi}_{k}+W_{k}R_{j}Z_{k}$
where $Z_{k}\sim\text{Dirichlet}(\delta)$ determines the direction of the deviation, $R_{j}\sim\text{Beta}(\omega J,J)$ controls the judge-specific magnitude, and $W_{k}\sim\text{Beta}(\omega K,K)$ controls the candidate-specific magnitude. The hyperparameter $\omega\in[0,\infty)$ controls the degree of constancy violation through the random effect magnitude, where $\omega=0$ enforces perfect constancy and increasing $\omega$ allows progressively larger violations. For instance, $\omega=1$ expects random effects to contribute half of the effect, while $\omega=8$ expects a contribution of 90%.
5.3 Prior for varying judge quality
Figure 5: Weight propagation framework for encoding judge quality priors in 3-level scoring, where outgoing edge weights must sum to 1. The probability of the judge assigning score $m_{2}$ when the true score is $m_{1}$, $\theta_{m_{1},m_{2}}$, equals the weighted sum of all its parent nodes per the incoming edge weights.
To construct a Bayesian prior over slopes between judge vertices, we parameterize the judge vertices in terms of a weight propagation graph (Figure B.10), where each element in the confusion matrix is an average of its parent nodes weighted by edge weights $\alpha$, i.e., $\theta_{\texttt{v}}^{(j)}=\sum_{\texttt{u}\rightarrow\texttt{v}}\theta_{\texttt{u}}^{(j)}\alpha_{\texttt{u}\rightarrow\texttt{v}}$. As such, the slopes between vertices are directly controlled in terms of edge weights. For each node v, we then place a Dirichlet prior over the outgoing edge weights $\alpha_{\texttt{v}}$ indexed by a single hyperparameter $\beta_{\max}$, where increasing $\beta_{\max}$ corresponds to more discriminative judges. For instance, in 3-level scoring, the Dirichlet prior for outgoing edges from node $\theta_{m_{1},m_{2}}$ has parameters $\vec{\beta}_{m_{1},m_{2}}=[1,1+\rho\cdot\beta_{\max},1]$ with $\rho\sim\text{Beta}(1,1)$. Full details are in the Appendix.
5.4 Conducting sensitivity analyses
Using this framework, we can develop a practical framework for conducting sensitivity analyses, to understand how epistemic uncertainty influences our confidence in the estimated rankings. In the absence of strong prior knowledge about LLM judge behavior, we recommend the following:
  1. 1.
    Start with conducting Bayesian inference with hyperparameters $\omega=0$ (perfect constancy) and $\beta_{\max}=5$ (moderate judge quality).
  2. 2.
    Estimate how rankings shift as the constancy assumption is relaxed, by increasing $\omega$ from 0 to 8 while holding $\beta_{\max}$ fixed.
  3. 3.
    Estimate how rankings shift as our prior over judge quality varies, by varying $\beta_{\max}$ from 0 to 20 while holding $\omega=0$.
If rankings remain stable across parameter variations, the conclusions are reliable despite epistemic uncertainty. Otherwise, one may need to inject additional prior knowledge (e.g., narrow the range of plausible hyperparameter values) or obtain gold-standard labels.
6 Experiments
Figure 6: Top: Sensitivity of estimated rankings when varying random effects (RE, left) and judge quality (right) hyperparameters, by plotting the correlation between the estimated ranking for each hyperparameter and their base values ($\omega=0,\beta_{\max}=0)$. Bottom: Candidates visualized on the probability simplex with judge configurations sampled from posterior (blue triangles), for MTBench (left) and Omni-MATH (right). We collapse 5-level ratings to 3 levels for visualization (see Appendix for mapping details).
Here we present experiments across various LLM benchmarks to assess the impact of epistemic uncertainty on ranking. First, using the geometric Bayesian framework, we conduct sensitivity analyses to determine the robustness of learned rankings. Second, we compare the overall performance of the Bayesian framework to existing methods in terms of coverage and correlation with ground truth rankings. We present our main experimental results below, with ablation studies, implementation details, and extended analyses in the Appendix.
We evaluate the Bayesian framework across three categories of benchmark datasets: (i) Verifiable tasks with 2-level scores: GPQA (Rein2023-dv) and MMLU Pro (Wang2024-ic) contain multiple-choice questions graded as correct/incorrect, with judges having an abstention option. (ii) Multi-level human-judged tasks: MTBench (Zheng2023-wr) evaluates multi-turn conversations while SummEval (Volske2017-bv) assesses summarizations. Both datasets are assessed on multi-level Likert scales, by LLMs as well as human experts. (iii) “Semi-verifiable” tasks: Omni-MATH (Gao2024-ec) contains mathematical reasoning problems with reference solutions, though there is no single ground-truth answer. LLM judges are asked to evaluate answers on a 3-level scale (correct, partial credit, incorrect).
We use a two-stage protocol to compare ranking methods. First, LLM judges evaluate candidates without access to ground truth, mirroring real-world usage. Second, we generate ground-truth scores for each answer by comparing against the correct multiple-choice answer on verifiable tasks and obtaining human-assigned scores for human-judged tasks. For semi-verifiable tasks, we rescore each candidate’s answer by providing the LLM judge with the provided reference answer. For GPQA, MMLU Pro, and Omni-MATH, we evaluated 19 contemporary models including Claude, GPT, Gemini, Llama, Mistral, and Qwen variants, with Claude 3.5 Haiku and GPT-4o mini serving as judges. These judges were deliberately selected to assess the performance of ranking methods when the judges are imperfect. For MTBench and SummEval, we utilize their existing candidates and judges.
6.1 Sensitivity analyses
Figure 7: Estimated rankings for the top 10 candidates on GPQA, presented in order by their true rankings. Unlike existing methods, Bayesian method fully incorporates epistemic uncertainty.
Table 1: Ranking method performance on benchmarks in terms of correlation and coverage of ground-truth ranks.
Method GPQA MMLU Omni-MATH SummEval MTBench Corr Cov Corr Cov Corr Cov Corr Cov Corr Cov Bayesian 0.881 (0.79, 0.90) 0.778 (0.68, 0.93) 0.862 (0.79, 0.88) 0.947 (0.95, 1.00) 0.696 (0.56, 0.75) 0.526 (0.38, 0.74) 0.888 (0.76, 0.91) 0.917 (0.67, 0.95) 1.000 (0.94, 1.00) 1.000 (1.00, 1.00) Bootstrap 0.847 (0.77, 0.89) 0.500 (0.56, 0.78) 0.826 (0.78, 0.87) 0.421 (0.47, 0.68) 0.682 (0.64, 0.69) 0.316 (0.26, 0.42) 0.885 (0.78, 0.90) 0.729 (0.65, 0.96) 0.943 (0.94, 1.00) 1.000 (1.00, 1.00) Bradley-Terry 0.847 (0.80, 0.88) 0.444 (0.40, 0.65) 0.842 (0.81, 0.87) 0.316 (0.33, 0.57) 0.682 (0.63, 0.69) 0.263 (0.21, 0.32) 0.902 (0.82, 0.91) 0.771 (0.61, 0.91) 0.943 (0.94, 1.00) 1.000 (0.67, 1.00) Simple Average 0.847 (0.77, 0.89) 0.167 (0.01, 0.28) 0.826 (0.78, 0.87) 0.105 (0.11, 0.21) 0.682 (0.64, 0.69) 0.158 (0.11, 0.21) 0.885 (0.78, 0.90) 0.271 (0.09, 0.54) 0.943 (0.94, 1.00) 0.667 (0.67, 1.00) Single Judge 0.817 (0.72, 0.87) 0.167 (0.01, 0.32) 0.825 (0.78, 0.87) 0.184 (0.11, 0.24) 0.632 (0.61, 0.67) 0.158 (0.11, 0.24) 0.787 (0.58, 0.86) 0.285 (0.06, 0.45) 0.971 (0.94, 1.00) 0.833 (0.67, 1.00)
For each benchmark dataset, we conducted Bayesian inference with random effects hyperparameter $\omega$ values of 0, 1, 2, 4, and 8, corresponding to expected random effects magnitudes of 0, 0.5, 0.7, 0.8 and 0.9 respectively. Figure 6 (left) shows that rankings remain remarkably stable across these variations for most datasets. The correlation between rankings at different $\omega$ values and the baseline rankings exceeds 0.95 for GPQA, MTBench, and SummEval, indicating that significant violations of the constancy assumption have minimal impact on the resulting rankings. Similarly, varying the judge quality hyperparameter $\beta_{\max}$ from 1 to 20 produces negligible effects on rankings across all datasets (Figure 6 right), with correlations consistently above 0.99. However, both MMLU Pro and Omni-MATH show greater sensitivity to the constancy assumption, with correlations between baseline rankings dropping to 0.77 and 0.86 respectively when $\omega$ is increased from 0 to 8. This sensitivity appears pronounced for MMLU Pro for $\omega>4$, highlighting uncertainty in its rankings because of potential variability in judge performance across candidates.
To understand why estimated rankings are more or less robust, we can visualize the candidates in the probability simplex against possible judge configurations sampled from the posterior distribution of the Bayesian method. Here we compare two contrasting datasets to illustrate why some benchmarks are robust to epistemic uncertainty while others are not. As shown in Figure 6 left, estimated rankings for MTBench are robust even when the constancy assumption is relaxed because (i) candidates are well-separated across the simplex, with clear performance tiers and (ii) the posterior judge configurations are relatively consistent with similar triangular shapes with minimal variability. In contrast, Figure 6 right shows that estimated rankings for Omni-MATH are more sensitive to epistemic uncertainty because (i) candidates cluster tightly in the central region with overlapping score distributions, making their relative ordering sensitive to judge vertex positions, and (ii) the posterior samples demonstrate considerable uncertainty in the judge configurations.
6.2 Comparative performance evaluation
We now evaluate the overall performance of the Bayesian framework against established baselines. To summarize the performance of the Bayesian procedure in a single metric, we fully incorporate epistemic uncertainty during posterior inference by defining a hyperprior over the hyperparameters varying the degree of constancy violation and judge quality. We compare against single judge scoring, score averaging across multiple judges (i.e., LLM judges as a jury), bootstrap confidence intervals (Xie2009-wv), and Bradley-Terry pairwise comparison models (Bradley1952-kj). Performance was evaluated in terms of Spearman correlation with the ground truth rankings and coverage rates of 95% credible/confidence intervals, averaged across candidates.
The Bayesian framework consistently achieves higher correlations and better coverage across all benchmark datasets (Table 1), with coverage rates improving by over 20 percentage points in some cases. This dramatic improvement in coverage reflects our theoretical finding that epistemic uncertainty is essential for proper uncertainty quantification. The correlation improvements, while more modest, demonstrate that incorporating judge uncertainty also benefits ranking estimates. Figure 7 illustrates how this coverage improvement is achieved: as seen on the GPQA dataset, the intervals from Bayesian inference are indeed significantly wider than those produced by comparator methods. Posterior distributions for the other datasets are given in the Appendix.
7 Discussion
This work leverages a novel geometric framework for understanding when gold-standard rankings can and cannot be recovered when one has access to imperfect LLM judges but no gold-standard labels. By representing judges and candidates on a probability simplex, we establish theoretical conditions for ranking identifiability and use these results to develop a sensitivity analysis approach. Our key insight is that epistemic uncertainty—uncertainty about judge quality—critically affects ranking reliability and is not considered by existing methods. The geometric perspective reveals that identifiability depends on scoring levels, judge consistency, and candidate separation in the simplex.
Appendix A Geometric Perspective: Additional details
A.1 Expected scores map to height in augmented space
A key geometric insight used throughout our analysis is that a candidate’s expected score corresponds to its height when the probability simplex is embedded in a higher-dimensional space. This result connects the abstract barycentric coordinates to the concrete ranking task.
Theorem 3 (Height Correspondence in Augmented Simplex).
Consider embedding the (M1)(M-1)-dimensional probability simplex into $M$-dimensional space by adding a “score” axis. Lift each judge vertex $\theta_{m}^{(j)}$ to position $(\theta_{m}^{(j)},m)$, placing it at height $m$. For any candidate $k$ with barycentric coordinates $\vec{\pi}_{k}=(\pi_{k,1},\ldots,\pi_{k,M})$, when positioned in the augmented space using these same coordinates, its height equals its expected score $\mathbb{E}[S_{k}^{*}]$.
Proof.
The candidate’s position in the original (M1)(M-1)-dimensional simplex is determined by its barycentric coordinates:
$\gamma_{k}^{(j)}=\sum_{m=1}^{M}\pi_{k,m}\theta_{m}^{(j)}$ (2)
where $\pi_{k,m}$ represents the weight given to judge vertex $\theta_{m}^{(j)}$.
To augment the space, we embed the simplex into $M$-dimensional space by adding a “score” coordinate. Each judge vertex $\theta_{m}^{(j)}$ is lifted to position $(\theta_{m}^{(j)},m)$, placing it at height equal to its score value $m$.
Crucially, when we position the candidate in this augmented space using the same barycentric coordinates $\pi_{k,m}$, we get:
$\text{Augmented position}=\sum_{m=1}^{M}\pi_{k,m}\left(\theta_{m}^{(j)},m\right)$ (3)
This convex combination yields:
  • Simplex coordinates: $\gamma_{k}^{(j)}=\sum_{m=1}^{M}\pi_{k,m}\theta_{m}^{(j)}$ (unchanged)
  • Height coordinate: $h_{k}=\sum_{m=1}^{M}\pi_{k,m}\cdot m$
The height $h_{k}=\sum_{m=1}^{M}\pi_{k,m}\cdot m$ is precisely the definition of the expected score $\mathbb{E}[S_{k}^{*}]$, where $\pi_{k,m}=\Pr(S_{k}^{*}=m)$.
Thus, the candidate’s height in the augmented space directly encodes its expected score, providing a visual representation where ranking candidates corresponds to comparing their heights. ∎
This geometric correspondence provides intuitive visualization: candidates automatically acquire heights equal to their expected scores when the simplex is embedded in augmented space using the barycentric coordinates. Ranking becomes a simple matter of comparing heights. Moreover, this representation naturally captures how uncertainty in judge vertex positions (epistemic uncertainty) translates to uncertainty in candidate heights and thus rankings.
A.2 Recoverability is dataset-specific
Figure A.8: Example where candidate rankings are likely identifiable even when there are 3+ levels, due to characteristics of the dataset. Candidates have assigned scores that are very different (mostly 1’s versus mostly 3’s).
There exist special cases where rankings become robust to judge quality assumptions despite the general non-identifiability result for 3+-level scoring systems. When candidates exhibit substantially different score distributions—manifesting as widely separated points on the probability simplex—their relative ordering becomes stable across a broad range of plausible judge configurations. See Figure A.8 as an example (Example 2 in Section 4.3).
Appendix B Bayesian model
B.1 Probability Model
For judge $j$ evaluating candidate $k$’s answer to the $i$-th question, we assume the assigned score $\hat{S}_{ik}^{(j)}$ given its true score $S_{ik}^{*}$ follows a multinomial distribution with parameter $\theta_{S_{ik}^{*},k}^{(j)}$ (Figure B.9). To maintain computational tractability while preserving the geometric structure identified in Section 3, we assume conditional independence across judges and questions. After marginalizing over the true latent scores, the likelihood of the observed data becomes:
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big],$ (4)
where $n$ is the number of questions and we exclude self-evaluations ($k\neq j$) to avoid self-preference bias.
Figure B.9: Plate diagram of the Bayesian model